HUBBLE SPACE TELESCOPE DAILY REPORT #2337 PERIOD COVERED: 0000Z (UTC) 03/25/99 - 0000Z (UTC) 03/26/99 Daily Status Report as of 085/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.2 Completed FOC/96 6350 (FOC spectroscopy of BALQSOs) The Faint Object Camera (f/96) was used to make ultraviolet spectrographic images of BALQSO 1331-0108 which is known to show broad emission lines from abundant ions. This will help us improve our knowledge of these BAL class objects. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7469 (Internal Velocity Distribution in Globular Clusters) The WF/PC-2 was used to make observations of the globular cluster NGC-104-R1 in order to study the internal velocity distribution in its cores and as a function of distance from the cluster center, and thus also infer its mass distribution. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The second iteration of this proposal was affected completely by the acquisition failure discussed in 2.1 and HSTAR 7073. The take data flag remained down and one abservation was lost. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.6 Completed STIS/CCD 8028 (STIS GTO Observations) The Space Telescope Imaging Spectrograph (CCD) was used to make GTO of stars in galaxy NGC-3369. The solar blind and long slit capabilities of STIS were used to investigate the population of hot stars in this galaxy. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD 7403 (Measuring Black Hole Masses in Broad-Lined AGNs) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of broad-lined active galactic nuclei (AGN) located in galaxy NGC-4036. Broad emission lines in AGNs are considered one of the most direct signatures of a large central mass, which is presumed to be a black hole. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 6701 (Photometry of the Globular Cluster System of the Sagittarius Dwarf Galaxy) The WF/PC-2 was used to make observations of the globular cluster system TER8 which is located the Sagittarius Dwarf Galaxy. Isochrone fits to these data will yield the accurate age of this system. The acquisition failure described in 2.1 and HSTAR 7073 affected approximately the first hour of this four-hour proposal. While the take data flag was down, five observations were lost. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 3 As described in HSTAR 7073, following two failed full mode FHST updates at 084/232307Z and at 084/232552Z, the acquisition at 084/232858Z also failed when the search radius limit was exceeded. An ARU/PRT was uplimked at 085/003217Z and the re-acquisiton at 085/011138Z was successful. The proposals described in 1.5 and 1.8 were affected. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 12 Successful : 10 See 2.1 for details of the two update failures. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared once. The STIS MCE-1 reset at 084/1822Z while the high voltage was off and while in an SAA interval. The STIS flight software error counter was cleared using ROP NS-12. The MAMA-1 was recovery via nominal SMS commanding at 085/0503Z. The NICMOS FPA temperature sensor for detector 1 was enabled at 084/2145Z per an operations request. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH